A job order cost sheet for Ryan Company is shown below Job No. 92 For 2,000 Units Direct Materials Labor Direct Manufacturing Date Overhead Beg. bal. 5,9007,400 5,180 Jan. 1 7,400 12 25 27 6,800 5,100 2,400 4,400 15,700 18,600 3,300 13,580 Cost of completed job: Direct materials 15,700 18,600 13,580 $47,880 $23.94 Direct labor Manufacturing overhead Total cost Unit cost ($47,880 2,000) (a) On the basis of the foregoing data, answer the following questions. (1) What was the balance in Work in Process Inventory on January 1 if this was the only unfinished job? Balance in Work in Process Inventory on January 1 18480 (2) If manufacturing overhead is applied on the basis of direct labor cost, what overhead rate was used in each year? (Round answers to o decimal places, eg, 55%.) Last year Crrent year Overhead rate (b) Prepare summary entries at January 31 to record the current years transactions pertaining to Job No. 92. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)

(b) Prepare summary entries at January 31 to record the current years transactions pertaining to Job No. 92. (Credit account titles are automatically indented when amount is entered. Do not indent manually.) Date Account Titles and Explanation Jan. 31 Work in Process Inventory Debit Credit 47880 Raw Materials Inventory 47880 To record raw materials used) 31 Work in Process Inventory 47880 Factory Labor To record factory labor used) 31 Work in Process Inventory 47880 Manufacturing Overhead 47880 To record manufacturing overhead) 31 Finished Goods Inventory 47880 Work in Process Inventory 47880 To record job completed)
Answer #1

a) 1. Balance in Work in Progress on January 1 = Direct Materials + Direct Labor + Manufacturing Overheads

= 5900+7400+5180 = 18480

2. Overhead rate = Manufacturing Overhead / Direct Labor

Last year = 5180 / 7400 = 70%

Current year = 5100/ 6800 = 75%

b)

Date Account Titles and Explanation Debit Credit
Jan.31 Work in Process Inventory 9800
Raw Materials Inventory 9800
[7400+2400]
(To record raw materials used)
Work in Process Inventory 11200
Factory Labor 11200
[6800+4400]
(To record factory labor used)
Work in Process Inventory 8400
Manufacturing overhead 8400
[5100+3300]
(To record manufacturing overhead)
Finished Goods Inventory 47800
Work in Process Inventory 47800
(To record job completed)